The turnbuckle is tightened until the tension in the cable AB equals 2.5 kN. Determine the vector expression for the tension T as a force acting on member AD. Also find the magnitude of the projection of T along the line AC. 5.0 m \1.4 n 2.1 m、y 3.5 m Answers: k) KN TAC = KN
Answer #1

Find the position vectors:

$$ \begin{array}{l} \mathrm{AB}=2.1 \mathrm{i}+1.4 \mathrm{j}-5 \mathrm{k} \mathrm{m} \\ \mathrm{AC}=2.1 \mathrm{i}-2.1 \mathrm{j}-5 \mathrm{k} \quad \mathrm{m} \end{array} $$

Find the unit vectors:

$$ \begin{array}{l} \lambda_{\mathrm{AB}}=\frac{2.1 \mathrm{i}+1.4 \mathrm{i}-5 \mathrm{k}}{\sqrt{2.1^{2}+1.4^{2}+(-5)^{2}}} \\ \lambda_{\mathrm{AC}}=\frac{2.1 \mathrm{i}-2.1 \mathrm{j}-5 \mathrm{k}}{\sqrt{2.1^{2}+(-2.1)^{2}+(-5)^{2}}} \end{array} $$

Find thetension vector:

$$ \mathbf{T}=T \lambda_{\mathrm{AB}}=2.5 \frac{2.1 \mathrm{i}+1.4 \mathrm{i}-5 \mathrm{k}}{\sqrt{2.1^{2}+1.4^{2}+(-5)^{2}}} \cdot=0.937 \mathrm{i}+0.625 \mathrm{j}-2.33 \mathrm{k} \mathrm{kN} $$

Find the projection of the \(\mathrm{T}\) along \(A C\) :

$$ \begin{aligned} \text { Projection } T_{A C} &=\mathbf{T} \cdot \lambda_{\mathrm{AC}} \\ &=(0.937 \mathrm{i}+0.625 \mathrm{j}-2.33 \mathrm{k}) \cdot\left(\frac{2.1 \mathrm{i}-2.1 \mathrm{j}-5 \mathrm{k}}{\sqrt{2.1^{2}+(-2.1)^{2}+(-5)^{2}}}\right) \\ &=2.11 \mathrm{kN} \end{aligned} $$